# Bulk-Boundary Correspondence and Exceptional Points for a Dimerized Hatano-Nelson Model with Staggered Potentials

### v1

@@ Abstract
It is well-known that the standard bulk-boundary correspondence does not hold for non-Hermitian systems in which also new phenomena such as exceptional points do occur. Here we study by analytical and numerical means a paradigmatic one-dimensional non-Hermitian model with dimerization, asymmetric hopping, and imaginary staggered potentials. We present analytical solutions for the eigenspectrum of this model with both open and closed boundary conditions as well as for the singular-value spectrum. We explicitly demonstrate the proper bulk-boundary correspondence between topological winding numbers in the periodic case and singular values in the open case. We also show that a non-trivial topology leads to protected eigenvalues in the entanglement spectrum. In the $\mathcal{PT}$-symmetric case, we find that the model has a phase where exceptional points become dense in the thermodynamic limit.
@@ I Introduction
It is well known that fundamental principles in quantum mechanics are formulated using Hermitian operators, which assure for example unitary time evolution of closed systems. However, practically, no quantum system is perfectly isolated. The way a system interacts with its environment is often represented through a Master equation like the Lindblad equation [1]. If quantum jumps can be neglected, then the time evolution described by a Lindblad equation can be replaced by an effective non-Hermitian (NH) Hamiltonian [2]. The validity of this effective approach has been verified in different experimental arrangements [3, 4, 5]. Strictly speaking, it requires a post-selection of those time evolutions which stay in the considered manifold [6].
@@
Non-Hermitian systems exhibit distinct properties associated with their complex spectrum and the difference between left and right eigenvectors [7]. In this context, as with Hermitian systems, Gaussian non-Hermitian models can offer valuable insights into possible phases and phenomena. In particular, the topological properties of Gaussian non-Hermitian models have been studied and new phenomena have been revealed [8, 9, 10, 11, 12, 13, 11, 14, 15, 16]. For example, non-hermiticity leads to a proliferation of possible symmetry classes: the famous Altland-Zirnbauer ten-fold classification [17] for Hermitian Gaussian models is replaced by a 38-fold classification [8, 14] in the non-Hermitian domain. The richer classification scheme is a direct consequence of the distinction between transposition and conjugate transposition. Other distinctive properties include the extreme sensitivity of the eigenspectrum to the boundary conditions [15] leading to the skin effect and the breakdown of the conventional bulk-boundary correspondence [10, 12, 18, 19]. Another remarkable aspect of non-Hermitian Hamiltonians is the existence of so-called exceptional points [20] where both eigenvalues and eigenvectors coalesce. Such exceptional points can be detected in experiments [6].
@@
In order to highlight the different phenomena possible in Gaussian non-Hermitian models, we will study a minimal model which allows to turn various symmetries on and off. In the Hermitian case, one of the best-known one-dimensional systems with symmetry-protected topological order is the Su-Schrieffer-Heeger (SSH) model [21] which has chiral symmetry and belongs to the BDI class. This model has alternating strong and weak hoppings and topological edge modes are present if the two sites at the end of an open chain are weakly coupled to their respective neighbors. Here we make the model non-Hermitian by allowing for parity breaking hoppings, i.e., different hopping amplitudes to the left and to the right. In addition, we also introduce staggered imaginary onsite potentials to allow for a non-Hermitian, $\mathcal{PT}$-symmetric limit. Physically, the parity breaking can arise in an effective description of an SSH model with gain and loss. This phenomenon has been investigated in plasmonic chains [22] and photonic lattices [23]. For equal hoppings to the left and to the right, the model is $\mathcal{PT}$ symmetric. The model is depicted in Fig. 1
@@
Figure 1: The considered minimal model has a two-site unit cell, asymmetric intra- ($V_{L},V_{R}$) and inter-cell ($W_{L},W_{R}$) hopping as well as complex staggered potentials ±$\pm iu$.
@@
Our paper is organized as follows: In Sec. II, we introduce the model and the difference equations the right and left eigenvectors have to fulfill. In Sec. III, we diagonalize the model for periodic and aperiodic boundary conditions. Furthermore, we determine the parameters for which the excitation gap closes and for which exceptional points do occur. We also introduce and determine the two winding numbers for this system. In Sec. IV, we then analytically study the open boundary case. As expected, the spectrum and the eigenvectors change completely as compared to the closed boundary case. This is also true for the conditions for the closing of the excitation gap and for the location of exceptional points. In Sec. V, we present some of the central results of our paper. We analytically calculate the singular value spectrum of a finite open chain as well as the zero-energy eigenstates of a semi-infinite chain. Based on these results we explicitly establish a bulk-boundary correspondence for this specific model following results recently proven in general in Ref. [25]. In Sec. VI, we then discuss the entanglement entropy when cutting the system in half. The studied model is Gaussian and therefore its correlation matrix can be obtained [26, 27, 28, 29, 30]. We find that in topological phases there are topologically protected eigenvalues in the entanglement spectrum. The last section is devoted to a short summary and our conclusions.
@@ II The dimerized Hatano-Nelson model
We consider a model with a two-site unit cell with asymmetric hopping in both intra ($V_{L},V_{R}$) and inter ($W_{L},W_{R}$) cells, see Fig. 1. Complex staggered fields ±$\pm\text{i}u$ act on the sites of the chain. This model can be viewed either as a dimerized Hatano-Nelson model or as a Su-Schrieffer-Heeger (SSH) model with asymmetric hopping. It is described by the Hamiltonian
@@
and $\bm{c}^{\dagger}=(c_{1}^{\dagger},\dots,c_{N}^{\dagger})$, $\bm{c}=(c_{1},\dots,c_{N})^{T}$ with $T$ denoting transposition.
To make this paper self-contained, we will recall the diagonalization of the asymmetric SSH chain for both closed ($\gamma=\pm 1)$ and open ($\gamma=0$) boundary conditions, referred to respectively as CBC or OBC. In both cases, the diagonalization of $\mathcal{T}_{N}$ with entries (II) follows from difference equations of the spectral problem. Here the lack of hermiticity of the single body Hamiltonian implies that left and right eigenvectors have to be considered. The spectral problems read
$\mathcal{T}_{N}\vec{r}=\epsilon~{}\vec{r},\quad(\mathcal{T}_{N})^{T}\vec{\ell}=\epsilon~{}\vec{\ell},$ (II.5)
where $\vec{x}=(x_{1},\dots,x_{N})^{T}$ are right ($x=r$) and left ($x=\ell$) eigenvectors of $\mathcal{T}_{N}$, while $\epsilon$ is an eigenvalue. The components of the right eigenvector satisfy the coupled difference equations
$V_{L}r_{2j}+W_{R}r_{2j-2}=(\epsilon-\text{i}u)r_{2j-1},$ (II.6)
$W_{L}r_{2j+1}+V_{R}r_{2j-1}=(\epsilon+\text{i}u)r_{2j},$ (II.7)
@@
$r_{0}=\gamma r_{N},\quad r_{N+1}=\gamma r_{1}.$ (II.8)
@@
$V_{R}\ell_{2j}+W_{L}\ell_{2j-2}=(\epsilon-\text{i}u)\ell_{2j-1},$ (II.9)
$W_{R}\ell_{2j+1}+V_{L}\ell_{2j-1}=(\epsilon+\text{i}u)\ell_{2j},$ (II.10)
@@
$\ell_{0}=\gamma\ell_{N},\quad\ell_{N+1}=\gamma\ell_{1}.$ (II.11)
For $\gamma=0$, we consider (II.8, II.11) for both parities of $N$, while for $\gamma\neq 0$ we only consider the case where $N$ is even.
In the following sections, each boundary condition is considered separately.
@@ III.1 Diagonalization
Let us briefly recapitulate the diagonalization of $\mathcal{T}_{N}$ for closed boundary conditions (CBC), that is, $\gamma=\pm 1$. In this case, $\mathcal{T}_{N}$ is a block circulant matrix and can be easily diagonalized. We only consider $N$ even for CBC. The difference equations (II.6,II.7,II.8) and (II.9,II.10,II.11) are easily solved by the ansatz
@@
±$\langle\ell^{\pm}(k)|H(k)=\epsilon^{\pm}(k)\langle\ell^{\pm}(k)|,$ (III.6)
@@
±$\langle\ell^{\pm}(k)|$ $=$ $\frac{1}{\sqrt{N^{\pm}}}\left(\begin{array}[]{cc}a_{L}^{\pm}(k),&b_{L}^{\pm}(k)\end{array}\right),$ (III.15)
@@
To conclude the solution, the parameter $k$ is quantized by (II.8) and given by
@@ III.2 Gap closing and exceptional points
The closing of both the real and the imaginary parts is well captured by the non-Hermitian gap [31]
$\Delta=\text{min}_{k}|\Delta_{c}(k)|.$ (III.24)
The complex gap (III.23) closes when,
@@
For each value of $k$ quantized by (III.21,III.22), the condition (III.25) leads to constraints in the model parameters and can determine a rich variety of phase boundaries. Furthermore, it follows immediately from Eqs. (III.13,III.17) that at points where the complex gap (III.25) vanishes the eigenvectors $\{|r^{+}(k)\rangle,|r^{-}(k)\rangle\}$ and also $\{\langle\ell^{+}(k)|,\langle\ell^{-}(k)|\}$ coalesce. Gap closing points are therefore also always exceptional points. Exceptional points and their relation with the discriminant of the characteristic polynomial of the hopping matrix $\mathcal{T}_{N}$ are discussed further in App. B.
To analyze the manifold of exceptional points given by Eq. (III.25) it is convenient to first set $u=0$. In this case, the system has sublattice symmetry $\sigma^{z}H(k)\sigma^{z}=-H(k)$. [^1] In non-Hermitian systems, chiral and sublattice symmetries are not equivalent [14]. Indeed for $u=0$ the chiral symmetry $\sigma^{z}H(k)^{\dagger}\sigma^{z}=-H(k)$ is not satisfied since $\sigma^{z}H(k)^{\dagger}\sigma^{z}=-H(k)^{\dagger}$. Then, for every allowed quantized $k_{m}$ value and relaxing the requirement of real couplings, the solutions of $H_{1}=0$ or $H_{2}=0$ are
$\left(\frac{V_{L}}{W_{R}}\right)_{\text{EP}}=-e^{\text{i}k_{m}}\quad\text{or}\quad\left(\frac{W_{L}}{V_{R}}\right)_{\text{EP}}=-e^{\text{i}k_{m}}.$ (III.26)
where $k_{m}$ is given by (III.21,III.22). It follows that for most values of $k_{m}$ the coupling parameters which lead to a gap closing/exceptional point are complex. For finite $N$ and periodic boundaries, one can see that $m=N/2$ ($k=2\pi$) and $m=N/4$ ($k=\pi$, if $N/2$ is even) are the only possible values of $k$ that leads to real ratios,
@@
Fixing $V_{R}=W_{R}=1$ (no loss in generality if $u=0$), these exceptional lines can be visualized as a function of $V_{L}$ and $W_{L}$, see the left panel in Fig. 2. For finite $N$ and antiperiodic boundaries, the ratios (III.26) are real for $m=1/2(N/2-1)$ ($k=\pi$, if $N/2-1$ is even). In the thermodynamic limit, the unit circles are filled and real ratios occur at $k=\pi$ or $k=2\pi$ for both periodic and antiperiodic boundary conditions. We remark that the sublattice-symmetric model has been widely considered in the literature, see e.g. Refs. [11, 13, 32, 33, 34, 35]. Here we emphasize that the gap closing points and exceptional points coincide and that they can also occur in the complex parameter plane for every allowed, quantized value $k_{m}$.
Figure 2: The non-Hermitian gap $\Delta$ as a function of the model parameters for $N=100$ and periodic boundary conditions. In panels (a) and (b) we set $V_{R}=W_{R}=1$ and $u=0$ and $u=0.5$ respectively. In panel (c) we consider $V_{L}=V_{R}=V$, $W_{L}=W_{R}=W$ and the field $u=1$. The numbers in panel (a) indicate the two winding numbers $(\nu_{1},\nu_{2})$, the numbers in panel (b) the winding number $I_{1}$.
Now we consider a nonzero real field $u$. The field couples $H_{1}(k)$ and $H_{2}(k)$ and Eq. (III.25) no longer factorizes, giving a constraint for all model parameters for every possible value of $k$. Eq. (III.25) can be solved for any of the model parameters, leading in general to a complex number. As an example, in the central panel of Fig. 2, the non-Hermitian gap (III.24) is shown for fixed parameters $V_{R}=W_{R}=1$ and $u=0.5$ as a function of $V_{L}$ and $W_{L}$. In this case, we can observe the merging of the zero field phases $(\nu_{1},\nu_{2})=(0,0)$ and $(1,-1)$ (we discuss the meaning of these numbers below) into a single phase as well as the bending of the phases boundaries.
Another interesting case to consider is the $\mathcal{PT}$ symmetric case with $V_{L}=V_{R}=V$ and $W_{L}=W_{R}=W$ but the real field $u$ is arbitrary, such that $\sigma^{x}H(k)\sigma^{x}=H(k)^{*}$. In this case, the constraint (III.25) reduces to,
@@
To visualize the gap closing, we can set $u=1$ (no loss of generality) and plot the gap (III.24) as a function of $V$ and $W$, see the right panel of Fig. 2. The dark blue lines mark the the gap closing/ exceptional points, which only occur within the regions $|V-W|<1<|V+W|$ or $|V+W|<1<|V-W|$. In the thermodynamic limit, this region is gapless and filled with exceptional points. It thus constitutes an entire exceptional phase.
We end this subsection remarking that in addition to the collision of quasienergies $\Delta_{c}=0$, another type of degeneracy in the spectrum can also be found. This degeneracy occurs within each band ϵ±superscriptitalic-$\epsilon^{\pm}$ for different quasimomentum $k$. Namely, one can impose
$H_{1}(k)H_{2}(k)=H_{1}(k^{\prime})H_{2}(k^{\prime})$ (III.29)
which gives another constraint on the model parameters. However, these degeneracies do not lead to exceptional points, as the eigenvectors do not coalesce, recall (III.13,III.17). This type of degeneracy, however, does appear as a factor in the discriminant of the characteristic polynomial of $\mathcal{T}_{N}$, see further remarks on this point in App. B, in particular Eq. (B.3).

### v3

# Exceptional Points, Bulk-Boundary Correspondence, and Entanglement Properties for a Dimerized Hatano-Nelson Model with Staggered Potentials

@@ Abstract
It is well-known that the standard bulk-boundary correspondence does not hold for non-Hermitian systems in which also new phenomena such as exceptional points do occur. Here we study, mostly by analytical means, a paradigmatic one-dimensional non-Hermitian model with dimerization, asymmetric hopping, and imaginary staggered potentials. We present analytical solutions for the singular-value and the eigensystem of this model with both open and closed boundary conditions. We explicitly demonstrate that the proper bulk-boundary correspondence is between topological winding numbers in the periodic case and singular values, not eigenvalues, in the open case. These protected singular values are connected to hidden edge modes which only become exact zero-energy eigenmodes in the semi-infinite chain limit. We also show that a non-trivial topology leads to protected eigenvalues in the entanglement spectrum. In the $\mathcal{PT}$-symmetric case, we find that the model has a so far overlooked phase where exceptional points become dense in the thermodynamic limit. This phase shows unusual hyper-ballistic transport properties with a dynamical critical exponent $z=1/2$.
@@ I Introduction
It is well known that fundamental principles in quantum mechanics are formulated using Hermitian operators, which assure for example unitary time evolution of closed systems. However, practically, no quantum system is perfectly isolated. The way a system interacts with its environment is often represented through a Master equation like the Lindblad equation [1]. If quantum jumps can be neglected, then the time evolution described by a Lindblad equation can be replaced by an effective non-Hermitian (NH) Hamiltonian [2]. The validity of this effective approach has been verified in different experimental arrangements [3, 4, 5]. Strictly speaking, it requires a post-selection of those time evolutions which stay in the considered manifold [6]. Another motivation to study non-Hermitian Hamiltonians came from realizing that hermiticity is not required to guarantee a real spectrum. The weaker condition of $\mathcal{PT}$-symmetry is sufficient, see e.g. Ref. [7] for a review.
@@
Generic non-Hermitian systems exhibit distinct properties associated with their complex spectrum and the difference between left and right eigenvectors [8]. In this context, as with Hermitian systems, Gaussian non-Hermitian models can offer valuable insights into possible phases and phenomena. In particular, the topological properties of Gaussian non-Hermitian models have been studied and new phenomena have been revealed [9, 10, 11, 12, 13, 14, 15, 12, 16, 17, 18]. For example, non-hermiticity leads to a proliferation of possible symmetry classes: the famous Altland-Zirnbauer ten-fold classification [19] for Hermitian Gaussian models is replaced by a 38-fold classification [9, 16] in the non-Hermitian domain. The richer classification scheme is a direct consequence of the distinction between transposition and conjugate transposition. Other distinctive properties include the extreme sensitivity of the eigenspectrum to the boundary conditions [13, 14, 17] leading to the skin effect and the breakdown of the conventional bulk-boundary correspondence [11, 14, 20, 21]. As has very recently been proven, the proper bulk-boundary correspondence in non-Hermitian systems is based on the theory of Toeplitz operators and is between winding numbers of the symbols of such operators in the periodic case and the singular-value spectrum in the open case [22]. Physically, it is important to note that the topologically protected singular values in general belong to boundary states which are metastable for finite systems and only become exact eigenstates in the limit of a semi-infinite system. The Hermitian case can then be understood as a special limit of this more general theory because in this case the singular values are just the absolute values of the eigenvalues. Another remarkable aspect of non-Hermitian Hamiltonians is the existence of so-called exceptional points [23] where both eigenvalues and eigenvectors coalesce. Such exceptional points can be detected in experiments [6].
@@
In order to highlight the different phenomena possible in Gaussian non-Hermitian models and to demonstrate in a fully analytical manner the recent novel results regarding their topology [22], we will study a minimal model which allows to turn various symmetries on and off. In the Hermitian case, one of the best-known one-dimensional systems with symmetry-protected topological order is the Su-Schrieffer-Heeger (SSH) model [24] which has chiral symmetry and belongs to the BDI class. This model has alternating strong and weak hoppings and topological edge modes are present if the two sites at the end of an open chain are weakly coupled to their respective neighbors. Here, we make the model non-Hermitian by allowing for parity breaking hoppings, i.e., different hopping amplitudes to the left and to the right. In addition, we also introduce staggered imaginary onsite potentials to allow for a non-Hermitian, $\mathcal{PT}$-symmetric limit. Physically, parity breaking can arise in an effective description of an SSH model with gain and loss. This phenomenon has been investigated in plasmonic chains [25] and photonic lattices [26]. The model is depicted in Fig. 1
@@
Figure 1: The considered minimal model has a two-site unit cell, asymmetric intra- ($V_{L},V_{R}$) and inter-cell ($W_{L},W_{R}$) hoppings as well as complex staggered potentials $\pm iu$.
@@
The main motivation for our paper is twofold: (1) To demonstrate a mathematically rigorous bulk-boundary correspondence using the apparatus developed in Ref. [22] for a concrete, physically relevant model. (2) To re-examine the phase diagram of a model which contains many limiting cases which have been studied previously in the literature but to do so, as far as possible, analytically and in a rigorous manner instead of using exact diagonalizations. In the process, we discover a novel exceptional phase with unusual properties.
Our paper is organized as follows: In Sec. II, we introduce the model and the difference equations the right and left eigenvectors have to fulfill. In Sec. III, we diagonalize the model for periodic and aperiodic boundary conditions. Furthermore, we determine the parameters for which the excitation gap closes and for which exceptional points do occur. We also point out that exceptional points can occur in the complex plane of the model parameters. We introduce and determine the two winding numbers for this system. In the $\mathcal{PT}$-symmetric phase we discover a phase where exceptional points are dense leading to hyper-ballistic transport. In Sec. IV, we then analytically study the open boundary case. As expected, the spectrum and the eigenvectors change completely as compared to the closed boundary case. This is also true for the conditions for the closing of the excitation gap and for the location of exceptional points. Depending on the parity of the length of the chain, we find different families of exceptional points which are determined via certain transcendental equations, also reflected in a factorization of the discriminant of the associated characteristic polynomial. They are clearly different from the closed boundary case. In Sec. V, we present some of the central results of our paper. We analytically calculate the singular value spectrum of a finite open chain as well as the zero-energy eigenstates of a semi-infinite chain. Based on these results we explicitly establish a bulk-boundary correspondence for this specific model following results recently proven in general in Ref. [22]. We explicitly show in particular that the topologically protected boundary modes are metastable for finite system size and only become exact eigenstates in the limit of a semi-infinite system. In Sec. VI, we then discuss in detail the entanglement entropy when cutting the system in half. The studied model is Gaussian and therefore its correlation matrix can be obtained [28, 29, 30, 31, 32, 33, 34, 35]. We find that in topological phases there are topologically protected eigenvalues in the entanglement spectrum. We note that the entanglement entropy is complex in general in the non-Hermitian case and find interesting properties of the entanglement entropy which can be directly tied to the physical properties of the system. In particular, we find that in the $\mathcal{PT}$-phase where exceptional points are dense the entanglement entropy is a nowhere continuous function. The last section is devoted to a short summary and our conclusions.
@@ II The dimerized Hatano-Nelson model
We consider a model with a two-site unit cell with asymmetric hopping in both intra ($V_{L},V_{R}$) and inter ($W_{L},W_{R}$) cells, see Fig. 1. Complex staggered fields $\pm\text{i}u$ act on the sites of the chain. This model can be viewed either as a dimerized Hatano-Nelson model or as a Su-Schrieffer-Heeger (SSH) model with asymmetric hopping. It is described by the Hamiltonian
@@
and $\bm{c}^{\dagger}=(c_{1}^{\dagger},\dots,c_{N}^{\dagger})$, $\bm{c}=(c_{1},\dots,c_{N})^{T}$ with $T$ denoting transposition. Note that in the semi-infinite case, the Hamilton operator has Toeplitz block form with $2\times 2$ matrices
$h_{0}=\begin{pmatrix}iu&V_{L}\\ V_{R}&-iu\end{pmatrix}\;,\;h_{1}=\begin{pmatrix}0&0\\ W_{L}&0\end{pmatrix}\;,\;h_{-1}=\begin{pmatrix}0&W_{R}\\ 0&0\end{pmatrix}$ (II.5)
which repeat along the diagonals. This will allow us later to define a corresponding symbol and winding numbers.
To make this paper self-contained, we will recall the diagonalization of the asymmetric SSH chain for both closed ($\gamma=\pm 1)$ and open ($\gamma=0$) boundary conditions, referred to respectively as CBC or OBC. In both cases, the diagonalization of $\mathcal{T}_{N}$ with entries (II) follows from difference equations of the spectral problem. Here the lack of hermiticity of the single-body Hamiltonian implies that left and right eigenvectors have to be considered. The spectral problems read
$\mathcal{T}_{N}~|\vec{r}\rangle=\epsilon~|\vec{r}\rangle,\quad\langle\vec{\ell}|~\mathcal{T}_{N}=\epsilon~\langle\vec{\ell}|,$ (II.6)
where $|\vec{r}\rangle=(r_{1},\dots,r_{N})^{T}$ and $\langle\vec{\ell}|=(\ell_{1},\dots,\ell_{N})$ are right and left eigenvectors of $\mathcal{T}_{N}$, respectively, with eigenvalue $\epsilon$.
The components of the right eigenvector satisfy the coupled difference equations
$V_{L}r_{2j}+W_{R}r_{2j-2}=(\epsilon-\text{i}u)r_{2j-1},$ (II.7)
$W_{L}r_{2j+1}+V_{R}r_{2j-1}=(\epsilon+\text{i}u)r_{2j},$ (II.8)
@@
$r_{0}=\gamma r_{N},\quad r_{N+1}=\gamma r_{1}.$ (II.9)
@@
$V_{R}\ell_{2j}+W_{L}\ell_{2j-2}=(\epsilon-\text{i}u)\ell_{2j-1},$ (II.10)
$W_{R}\ell_{2j+1}+V_{L}\ell_{2j-1}=(\epsilon+\text{i}u)\ell_{2j},$ (II.11)
@@
$\ell_{0}=\gamma\ell_{N},\quad\ell_{N+1}=\gamma\ell_{1}.$ (II.12)
For $\gamma=0$, we consider (II.9, II.12) for both parities of $N$, while for $\gamma\neq 0$ we only consider the case where $N$ is even. In the following sections, each boundary condition is considered separately.
@@ III.1 Diagonalization
Let us briefly recapitulate the diagonalization of $\mathcal{T}_{N}$ for closed boundary conditions (CBC), that is, $\gamma=\pm 1$. In this case, $\mathcal{T}_{N}$ is a block circulant matrix and can be easily diagonalized. We only consider $N$ even for CBC. The difference equations (II.7,II.8,II.9) and (II.10,II.11,II.12) are easily solved by the ansatz
@@
$\langle\ell^{\pm}(k)|H(k)=\epsilon^{\pm}(k)\langle\ell^{\pm}(k)|,$ (III.6)
@@
$\langle\ell^{\pm}(k)|$ $=$ $\frac{1}{\sqrt{N^{\pm}}}\left(\begin{array}[]{cc}a_{L}^{\pm}(k),&b_{L}^{\pm}(k)\end{array}\right),$ (III.15)
@@
To conclude the solution, the parameter $k$ is quantized by (II.9) and given by
@@ III.2 Gap closing and exceptional points
We define the non-Hermitian gap by [36]
$\Delta=\text{min}_{k}|\Delta_{c}(k)|$ (III.24)
which closes if
@@
For each value of $k$ quantized by (III.21,III.22), the condition (III.25) leads to constraints in the model parameters and can determine a rich variety of phase boundaries. Furthermore, it follows immediately from Eqs. (III.13,III.17) that at points where the complex gap (III.25) vanishes the eigenvectors $\{|r^{+}(k)\rangle,|r^{-}(k)\rangle\}$ and also $\{\langle\ell^{+}(k)|,\langle\ell^{-}(k)|\}$ coalesce. Gap closing points are therefore also always exceptional points. Exceptional points and their relation with the discriminant of the characteristic polynomial of the hopping matrix $\mathcal{T}_{N}$ are discussed further in App. A.
To analyze the phase transitions and the manifold of exceptional points given by Eq. (III.25) it is convenient to first set $u=0$. In this case, the system has sublattice symmetry $\sigma^{z}H(k)\sigma^{z}=-H(k)$. [^1] In non-Hermitian systems, chiral and sublattice symmetries are not equivalent [16]. Indeed for $u=0$ the chiral symmetry $\sigma^{z}H(k)^{\dagger}\sigma^{z}=-H(k)$ is not satisfied since $\sigma^{z}H(k)^{\dagger}\sigma^{z}=-H(k)^{\dagger}$. Then, for every allowed quantized $k_{m}$ value and relaxing the requirement of real couplings, the solutions of $H_{1}=0$ or $H_{2}=0$ are
$\left(\frac{V_{L}}{W_{R}}\right)_{\text{EP}}=-e^{\text{i}k_{m}}\quad\text{or}\quad\left(\frac{W_{L}}{V_{R}}\right)_{\text{EP}}=-e^{\text{i}k_{m}}$ (III.26)
where $k_{m}$ is given by (III.21,III.22). It follows that for most values of $k_{m}$ the coupling parameters which lead to a gap closing/exceptional point are complex. For finite $N$ and periodic boundaries, one can see that $m=N/2$ ($k=2\pi$) and $m=N/4$ ($k=\pi$, if $N/2$ is even) are the only possible values of $k$ that lead to real ratios,
@@
Fixing $V_{R}=W_{R}=1$ (no loss in generality if $u=0$), these exceptional lines can be visualized as a function of $V_{L}$ and $W_{L}$, see the left panel in Fig. 2. They separate gapped phases labeled by two integer winding numbers $(\nu_{1},\nu_{2})$ whose definition is discussed in subsection III.3. With respect to the type of gap [16], the phases (0,1)(0,-1) and $(1,0)$ have a point gap while phases $(0,0)$ and (1,1)(1,-1) have a line gap (real or imaginary). For finite $N$ and antiperiodic boundaries, the ratios (III.26) are real for m=1/2(N/21)m=1/2(N/2-1) ($k=\pi$, if N/21N/2-1 is even). In the thermodynamic limit, the unit circles are filled and real ratios occur at $k=\pi$ or $k=2\pi$ for both periodic and antiperiodic boundary conditions. We remark that sublattice-symmetric models have been widely considered in the literature, see e.g. Refs. [12, 15, 37, 38, 39, 40]. Here we emphasize that the gap closing points and exceptional points coincide and that they also occur in the complex parameter plane for every allowed quantized value $k_{m}$.
Figure 2: The non-Hermitian gap $\Delta$, see Eq. (III.24), as a function of the model parameters in the thermodynamic limit for periodic boundary conditions. In panel (a), the zero field (sublattice symmetric) case, $u=0$, is shown for $V_{R}=W_{R}=1$ with two winding numbers $(\nu_{1},\nu_{2})$ characterizing each phase. In panel (b), we set $u=0.5$, therefore breaking the sublattice symmetry and leaving only a single winding number $I_{1}$ labeling the emerging phases. In panel (c), we consider the $\mathcal{PT}$-symmetric case $V_{L}=V_{R}=V$, $W_{L}=W_{R}=W$ and $u=1$. The regions in blue constitute a gapless phase where exceptional points are dense. The central “diamond” is a complex (anti-$\mathcal{PT}$) phase. The remaining phases are $\mathcal{PT}$-unbroken phases, either trivial or topological.
Next, we consider a nonzero real field $u$. The field couples $H_{1}(k)$ and $H_{2}(k)$ and Eq. (III.25) no longer factorizes, giving a constraint for all model parameters for every possible value of $k$. Eq. (III.25) can be solved for any of the model parameters, leading in general to a complex number. As an example, in the central panel of Fig. 2, the non-Hermitian gap (III.24) is shown for fixed parameters $V_{R}=W_{R}=1$ and $u=0.5$ as a function of $V_{L}$ and $W_{L}$. The emerging phases are characterized by an integer winding number $I_{1}$ defined in subsection III.3. We can observe, for example, that the field $u$ induces the merging of the zero field phases $(\nu_{1},\nu_{2})=(0,0)$ and (1,1)(1,-1), with $V_{L}<-1$ and $W_{L}<-1$, into a single phase characterized by $I_{1}=0$. This can be understood by looking at the point $V_{L}=W_{L}=-1$, where $H_{1}(k)H_{2}(k)-u^{2}=-(2\sin(k/2)+\text{i}u)(2\sin(k/2)-\text{i}u)$ cannot be canceled for any real $k$ and $u$; therefore, the gap is open. On the other hand, the zero field phases $(0,0)$ and (1,1)(1,-1), with $V_{L}>1$ and $W_{L}>1$ do not merge at $u=0.5$, although the boundaries of the phases are bend. For instance, at $V_{L}=W_{L}=1$, one has $H_{1}(k)H_{2}(k)-u^{2}=(2\cos(k/2)-u)(2\cos(k/2)+u)$ which can be canceled by properly choosing a real $k$ depending on the value of the field, provided $|u|\leq 2$.
Another interesting case to consider is the $\mathcal{PT}$ symmetric case with $V_{L}=V_{R}=V$ and $W_{L}=W_{R}=W$ and $u$ real but arbitrary, such that $\sigma^{x}H(k)\sigma^{x}=H(k)^{*}$. In this case, the constraint (III.25) reduces to,
@@
To visualize the gap closing, we can set $u=1$ (no loss of generality) and plot the gap (III.24) as a function of $V$ and $W$, see the right panel of Fig. 2. The blue regions |VW|<1<|V+W||V-W|<1<|V+W| or |V+W|<1<|VW||V+W|<1<|V-W| are gapless in the thermodynamic limit. This is a so far overlooked phase where the exceptional points are dense. In this phase, the $\mathcal{PT}$-symmetry is broken. The regions |V||W|>1|V|-|W|>1 and |W||V|>1|W|-|V|>1 are characterized by a purely real spectrum, and are adiabatically connected to the standard SSH model via $u\rightarrow 0$. Then, the former is called the $\mathcal{PT}$-unbroken trivial phase and the latter the $\mathcal{PT}$-unbroken topological phase. In the phase $|W|+|V|<1$ the spectrum is purely imaginary, and we name it the complex phase or anti-$\mathcal{PT}$ phase.
The gapless phase where exceptional points are dense is peculiar because there is always a set of eigenvectors which is coalescent. I.e., this gapless phase is very different from a Hermitian phase where eigenvectors are orthogonal and one might wonder how this coalescence of eigenvectors affects its transport properties. To address this question, we investigate the scaling of the gap $\Delta\sim N^{-z}$ where $z$ is the dynamical critical exponent. A problem with such an analysis for a generic ratio of parameters $V/W$ is that the allowed momenta $4\pi m/N$ with $m=1,\cdots,N/2$ might or might not be commensurate or almost commensurate with the gap closing condition (III.28) depending on the system size $N$. In this case no clear scaling for generic $N$ emerges. To circumvent this issue, we choose a ratio $V/W$ which is badly approximable in the Diophantine sense by rational numbers thus making it possible to consider generic $N$. In Fig. 3 we show, in particular, results for $V/W$ being the golden or the inverse golden ratio, respectively.
Figure 3: Scaling of the gap (III.24) in the phase where exceptional points become dense for $V/W=(1-\sqrt{5})/2$ (black circles) and $V/W=(1+\sqrt{5})/2$ (red squares). The lines are fits $\Delta\sim N^{-1/2}$, i.e., the dynamical critical exponent is $z=1/2$.
The results indicate that the dynamical critical exponent is $z=1/2$ which means that lengths scale as $\ell\sim t^{2}$. Transport in this phase is thus hyper-ballistic. We note that in a short-range Hermitian lattice model the Lieb-Robinson bounds make faster than ballistic transport impossible. However, such bounds do not apply in a non-Hermitian model where eigenvectors do not need to be orthogonal. In this phase, in particular, eigenvectors coalesce leading to an amplification.